OpenStudy (anonymous):

factor and simplify the expression: cos^2x+cos^2xcot^2x

OpenStudy (anonymous):

hey, that's just cot^2(x)!

OpenStudy (amistre64):

cos^2 + cos^2 cos^2 ----------- is what I see so far sin^2

OpenStudy (anonymous):

=cos^2(x)(1+cos^2(x)/sin^2(x)) =cos^2(x)((sin^2(x)+cos^2(x))/sin^2(x)) =cos^2(x)/sin^2(x) =cot^2(x)

OpenStudy (anonymous):

work of art i say

OpenStudy (amistre64):

or factor out a cos^2 lol

OpenStudy (amistre64):

cos^2(1+cot^2) cos^2(csc^2) cos^2 ----- = 1 ...... cos^2

OpenStudy (amistre64):

let me check my 1+cot^2

OpenStudy (amistre64):

sin^2 + cos^2 = 1 ------------ ---- sin^2 sin^2 sin^2 1 + cot^2 = csc^2.... my bad lol

OpenStudy (amistre64):

cos^2 ----- = cot^2 smarty was right :) sin^2

OpenStudy (anonymous):

yours more elegant tho

OpenStudy (anonymous):

csc=1/sin^2

OpenStudy (anonymous):

i did it on the fly

OpenStudy (amistre64):

i did it with one synapse tied behind my right prefrontal lobe ;)